You’re a robotics engineer designing a hydraulic system to move a robotic arm. The hydraulic cylinder that drives the arm has diameter 5.0 cm and can exert a maximum force of 5.6 kN . Hydraulic tubing comes rated in multiples of 1/2 MPa, and for safety, you’re to specify tubing capable of withstanding 50% greater pressure than it will ever experience in use.

P=____ MPa
Answer #1

The Force per area = 5600N/π*0.025² = 2852056.58 N/m² = 2.85 MPa
1.5*2.85 = 4.275 MPa so you would order 4.275 MPa rated tubing because that is the nearest 1/2 MPa
Answer #2

SOLUTION 


Pressure of the system 

= Force / Area of piston 

= 5.6* 10^3 / (pi/4 * (5.0 * 10^(-2))^2) 

= 2.852 * 10^6 Pa

= 2.852 MPa .


Pipe should withstand pressure = 2.852 * 1.50 = 4.278 MPa.


Pipes are rated in the steps of 1/2 MPa,

So, pipe required is :  4 and 1/2 MPa  rating (ANSWER).

answered by: Tulsiram Garg